### v2

# Dynamical arrest in active nematic turbulence

Abstract
Active fluids display spontaneous turbulent-like flows known as active turbulence. Recent work revealed that these flows have universal features, independent of the material properties and of the presence of topological defects. However, the differences between defect-laden and defect-free active turbulence remain largely unexplored. Here, by means of large-scale numerical simulations, we show that defect-free active nematic turbulence can undergo dynamical arrest. We find that flow alignment—the tendency of nematics to reorient under shear—enhances large-scale jets in contractile rodlike systems while promoting arrested flow patterns in extensile systems. Our results reveal a mechanism of labyrinthine pattern formation produced by an emergent topology of nematic domain walls that partially suppresses chaotic flows. Taken together, our findings call for the experimental realization of defect-free active nematics, and suggest that topological defects enable turbulence by preventing dynamical arrest.
Active fluids are driven internally by their own components—be they molecular motors, cells, or synthetic colloidal particles. This internal driving produces spontaneous flows, which often exhibit spatio-temporal chaos at high activity 1. Chaotic flows have been observed in a variety of systems, such as bacterial suspensions 2; 3; 4; 5; 6; 7; 8; 9; 10; 11, sperm 12, mixtures of cytoskeletal components 13; 14; 15; 16; 17; 18; 19; 20; 21; 22, cell monolayers 23; 24; 25; 26; 27, and artificial self-propelled particles 28; 29; 30; 31. Although all these systems operate at low Reynolds number, for which inertia is negligible, the flows are reminiscent of classic inertial turbulence. Hence, active chaotic flows are generically referred to as active turbulence 1.
Active turbulence encompasses systems with different symmetries, including polar or nematic orientational order, each exhibiting unique statistical properties 1. Recent work showed that the velocity power spectrum of active nematic turbulence features scaling laws with universal exponents independent of the material properties of the fluid such as its viscosity and activity 32; 33; 22. Whereas most experiments and simulations prominently feature topological defects 34, turbulence with the same universal exponents was also found for strongly-ordered nematics in the absence of defects 33. However, how the material properties and the presence or absence of defects affect other, non-universal features of the flow field remains unclear.
Here, we show that defect-free active nematic turbulence exhibits arrested patterns not seen in the presence of defects. We reveal these patterns by studying the effect of the flow-alignment parameter $\nu$, which characterizes the tendency of liquid crystals to reorient under shear. As expected, we find that the scaling exponent of the velocity power spectrum is unaffected when varying $\nu$, confirming its universal character 33; 1. Despite keeping the same scaling, we find that flow-alignment strongly affects both the strength and the spatiotemporal structure of the flows.
We identify two distinct regimes depending on the parameter combination $\text{sign}(\zeta)\nu$, where $\zeta$ is the active stress parameter, with $\zeta>0$ for extensile and $\zeta<0$ for contractile stresses, respectively. We discuss the case of rod-like nematic components, for which $\nu\leq 0$. In the flow-aligning regime, where $\nu<-1$, we find that contractile nematics (with $\zeta\nu>0$) exhibit strongly chaotic flows and nematic distortions, which largely decorrelate from one another. In this regime, nematic domain walls—lines of strong variation of the director—appear, split, and dissolve, forming a dynamically reorganizing pattern (Figs. 1 to 1, Supplementary Movie 1). In contrast, for extensile aligning nematics ($\zeta\nu<0$), we find that domain walls are strongly stabilized by the flow they entrain. The walls typically grow and branch but neither break nor merge. As walls avoid each other, they become arrested (or grid-locked) in a space-filling tree-like pattern (Figs. 1 to 1, Supplementary Movie 2). Once arrested, the pattern exhibits slow residual dynamics that shares some features with ageing phenomena in glassy systems. The special case $\nu=0$ has been the focus of previous work33 and is considered here as a reference. The dynamics in this case lie between these two extremes while also manifesting comparatively shorter wavelengths (Supplementary Fig. 1, Supplementary Movie 3).
We uncover that the dynamical arrest stems from the emergence of an effective topology, characterized by a set of connectivity rules for the domain walls imposed by the combined effects of activity and flow alignment. In particular, we demonstrate that, in the absence of topological defects, the arrested state is organized by local nematic structures, which we name pseudo-defects. We discuss their topological interpretation and show that they coexist with actual defects in experiments on microtubule-based active nematics. While future studies could clarify which features of defect-free active flows may also be present in active nematics with defects, achieving defect-free active nematics remains a compelling experimental challenge.
Hydrodynamics of defect-free active nematics
We generalize our previous minimal model of active nematic turbulence 33 by adding flow alignment and the Ericksen stress 35. We work at vanishing Reynolds number, for which momentum balance reads:
$0=-\partial_{\alpha}P+\partial_{\beta}\left(\sigma_{\alpha\beta}^{\text{a}}+\sigma_{\alpha\beta}^{\text{E}}+\sigma_{\alpha\beta}\right).$ (1)
The pressure $P$ enforces the incompressibility condition $\partial_{\alpha}v_{\alpha}=0$ of the flow field $\bm{v}$. The next two terms are the divergence of the antisymmetric part of the stress tensor, and of the symmetric Ericksen stress, respectively35; 36:
$\sigma_{\alpha\beta}^{\text{a}}=\frac{1}{2}\left(n_{\alpha}h_{\beta}-h_{\alpha}n_{\beta}\right),\quad\sigma_{\alpha\beta}^{\text{E}}=-\frac{\delta F_{n}}{\delta\left(\partial_{\alpha}n_{\gamma}\right)}\partial_{\beta}n_{\gamma}.$ (2)
These stresses arise from the elastic distortions of the director field $\bm{n}$. Distortions produce an orientational field $h_{\alpha}=-\delta F_{n}/\delta n_{\alpha}=K\nabla^{2}n_{\alpha}+h_{\parallel}^{0}n_{\alpha}$ computed from the Frank free energy which, in the one-constant $K$ approximation, reads
[…]
Here, we assumed that the fluid is deep in the nematic phase so that the director field has a fixed modulus $|\bm{n}|=1$. This constraint is imposed by the second term through the Lagrange multiplier $h_{\parallel}^{0}$. We start with a continuous director field, which precludes the generation of topological defects 33; 37.
The last term in Eq. 1 arises from the symmetric part of the deviatoric stress, which is given by the constitutive equation 38; 39; 40; 36
$\sigma_{\alpha\beta}=2\eta v_{\alpha\beta}-\zeta\hat{q}_{\alpha\beta}+\frac{\nu}{2}\left(n_{\alpha}h_{\beta}+h_{\alpha}n_{\beta}-n_{\gamma}h_{\gamma}\delta_{\alpha\beta}\right).$ (4)
Here, $\eta$ is the shear viscosity, $v_{\alpha\beta}=1/2\left(\partial_{\alpha}v_{\beta}+\partial_{\beta}v_{\alpha}\right)$ is the symmetric part of the strain-rate tensor, $\zeta$ is the active stress parameter, and $\hat{q}_{\alpha\beta}=n_{\alpha}n_{\beta}-1/2\,\delta_{\alpha\beta}$ is the nematic orientation tensor defined by the director $\bm{n}$. The last term describes stresses due to flow alignment, with parameter $\nu$.
As in passive nematics, the dynamics of the director field are given by
$\partial_{t}n_{\alpha}+v_{\beta}\partial_{\beta}n_{\alpha}+\omega_{\alpha\beta}n_{\beta}=\frac{1}{\gamma}h_{\alpha}-\nu v_{\alpha\beta}n_{\beta},$ (5)
where $\omega_{\alpha\beta}=1/2\left(\partial_{\alpha}v_{\beta}-\partial_{\beta}v_{\alpha}\right)$ is the vorticity tensor, and $\gamma$ is the rotational viscosity. The left-hand side is the co-rotational material derivative of the director field. On the right-hand side, the orientational field $h_{\alpha}$ specifies the elastic torque acting on the director, whereas the last term captures its reorientation due to extensional flow, i.e., the flow-alignment effect.
Nondimensionalizing the equations, we rescale length by the system size $L$, time by the active time $\tau_{\text{a}}=\eta/|\zeta|$, pressure by the active stress $|\zeta|$, and orientational field by $K/L^{2}$. To eliminate the pressure, we take the curl of Eq. 1 and obtain a Poisson equation for the vorticity $\omega$, which we write in terms of the stream function $\psi$ (defined by $v_{x}=\partial_{y}\psi$ and $v_{y}=-\partial_{x}\psi$) 37, as $\omega=-\nabla^{2}\psi$:
$\nabla^{2}\omega=-\nabla^{4}\psi=-S\partial_{\alpha}\partial_{\beta}\bar{q}_{\alpha\beta}+R\nu\partial_{\alpha}\partial_{\beta}(\bar{q}_{\alpha\beta}h_{\parallel}-\hat{q}_{\alpha\beta}h_{\perp})\\ +\frac{R}{2}\nabla^{2}h_{\perp}+R\epsilon_{\alpha\beta}(\partial_{\alpha}h_{\parallel})\partial_{\beta}\theta.$ (6)
Here, $S$ is the sign of the active stresses, with $S=\pm 1$ for extensile and contractile stresses, respectively, and $\theta$ is the nematic director angle: $\bm{n}=(\cos\theta,\,\sin\theta)$. The right-hand side contains four sources of vorticity, arising respectively from active, flow-alignment, antisymmetric, and Ericksen stresses. We used the totally antisymmetric tensor $\epsilon_{\alpha\beta}$ to define $\bar{q}_{\alpha\beta}\equiv-\epsilon_{\alpha\gamma}\hat{q}_{\gamma\beta}$. We also introduced the parallel and perpendicular components of the orientational field 37:
$h_{\perp}=\frac{1}{A}\nabla^{2}\theta,\qquad h_{\parallel}=\nu\bar{q}_{\alpha\beta}\partial_{\alpha}\partial_{\beta}\psi.$ (7)
Equations 6 and 7 contain three dimensionless parameters: the activity number $A=L^{2}/\ell_{\text{a}}^{2}$, the viscosity ratio $R=\gamma/\eta$, and the flow-alignment parameter $\nu$. The activity number compares the system size $L$ with the active length $\ell_{\text{a}}=\sqrt{K/(|\zeta|R)}$ defined by the balance between active and elastic nematic stresses.
Finally, the director dynamics Eq. 5 breaks down into a component parallel to $\bm{n}$, which specifies $h_{\parallel}$ as given in Eq. 7, and a component perpendicular to $\bm{n}$, which determines the dynamics of the nematic angle $\theta$ 37:
$\partial_{t}\theta-\epsilon_{\alpha\beta}\partial_{\alpha}\psi\,\partial_{\beta}\theta+\frac{1}{2}\nabla^{2}\psi=h_{\perp}+\nu\hat{q}_{\alpha\beta}\partial_{\alpha}\partial_{\beta}\psi.$ (8)
Equations 6 to 8 give the hydrodynamics of our active fluid in terms two fields: the angle $\theta$ and the stream function $\psi$.
We numerically integrate Eqs. 6 to 8 in a square with periodic boundary conditions using a hybrid numerical scheme that combines a pseudo-spectral method for Eq. 6 with the finite element method for Eq. 8 (Methods .1). Unless otherwise specified, we set $R=1$ and $A=3.2\times 10^{5}$, for which the system was shown to develop large-scale active turbulence when $\nu=0$ 33. The dynamics are invariant under the transformation ($\nu\rightarrow-\nu$, $S\rightarrow-S$, $\theta\rightarrow\theta\pm\pi/2$), meaning that changing the signs of $\nu$ and $S$ simultaneously leaves the system invariant up to a rotation 41; 42; 37.
[…]
Figure 1: Strong and arrested regimes of active nematic turbulence. 11, Snapshots from simulations of defect-free active nematic turbulence in contractile (11) and extensile (11) flow-aligning systems. Parameter values were set to $R=1$, $\nu=-1.1$, and $A=3.2\times 10^{5}$. Top panels (1,1) show the flow field; black curves are streamlines, and the color indicates the speed (see Supplementary Movies 1 and 2). Middle panels (1,1) show the Frank free energy density $\sim|\nabla\theta|^{2}$, with high-intensity lines corresponding to nematic domain walls (see Supplementary Movies 1 and 2). Bottom panels (1,1) are zooms highlighting the type of nematic distortion as well as the interplay between nematic walls and flows. The gray-scale background is the line integral convolution representation of the director field $\bm{n}$. Magenta and cyan intensities respectively represent splay $(\nabla\cdot\bm{n})^{2}$ and bend |×$|\nabla\times\bm{n}|^{2}$ contributions to the Frank energy density. The black arrows represent the flow field $\bm{v}$, which localizes along the nematic walls in the arrested regime. Black circles indicate stagnation points of the flow. White scale bar represents the selected wavelength $\lambda_{i}$. 11, Spectra characterizing fully-developed active nematic turbulence (see details in Methods .2). The lines in (1,1) represent a smoothed (Gaussian) interpolation of the computed data points. We compare the flow-aligning contractile (red, as in 11 and Supplementary Movie 1) and extensile (blue, as in 11 and Supplementary Movie 2) cases with the $\nu=0$ case (black, as in Supplementary Fig. 1 and Supplementary Movie 3), for which contractile and extensile stresses are equivalent up to a rotation41; 42; 37 . 1, Velocity power spectrum on a log-log scale, showing (i) the universal low-$q$ scaling law and (ii) the distinct organization of flows across scales in the different cases. The wider scaling regime in the contractile case captures the strong large-scale jets (see 1). The peak in the extensile case is representative of wall streams (see 1). 1, Frank energy spectrum, showing that (i) the selected wavelength (peak position) depends on $\nu$ but not on the sign of active stress, and (ii) the peak width depends on the sign of active stress when $\nu\neq 0$. 1, Spectrum of correlation times associated with the flow $\bm{v}$ (light colored points and lines) and the nematic tensor $\hat{q}_{\alpha\beta}$ (darker points and lines). This log-log plot reveals strong differences in decay times between the regimes, as well as the differences between the flow and nematic tensor within a regime. Correlation times are extracted from exponential fits to the corresponding space-time autocorrelation functions in Fourier space (see Methods .2).
Strong large-scale turbulence
We first study the case $S\nu>1$, realized with contractile flow-aligning rods ($S=-1$ and $\nu=-1.1$). We find a chaotic flow field with vortices at many scales and strong (albeit transient) large-scale jets (Fig. 1). These flow patterns stem from long-range hydrodynamic interactions and are largely uncorrelated with the director field (Fig. 1), which also evolves chaotically. The Frank energy density concentrates along domain walls (Figs. 1 and 1)—well-known regions of strong orientation gradients 43; 44; 45. Such walls feature a typical separation length, corresponding to the peak position in the Frank energy spectrum (Fig. 1, red), coinciding with the peak of the nematic tensor power spectrum (Supplementary Fig. 2, red). Compared with the $\nu=0$ case33, the nematic pattern in the $S\nu>1$ regime appears more broken, with fragmented walls coexisting with wall-free regions (dark patches in Fig. 1). As a consequence, the total Frank free energy is smaller, and its peak is shifted to larger scales (Fig. 1, red). A dynamically-rearranging pattern emerges as walls spontaneously appear, split, and dissolve (Supplementary Movies 1 and 4, left). Both the flows and the nematic tensor are very dynamic, as confirmed by their short correlation time at all length scales (Fig. 1, red).
In recent work we have shown that, for $S\nu>1$, both the quiescent state with uniform orientation and the flowing state with splaying domain walls are non-linearly unstable against one-dimensional modulations37. Thus, strong enough perturbations can trigger sudden transitions between these states. For this reason, we assert that domain wall patterns can be disrupted by the strong turbulent flows. In the two-dimensional simulations shown here, we observe the coexistence of uniform and wall-laden patches that continuously rearrange. We conclude that the metastibility associated with this regime strongly enhances the chaotic flows, particularly at large-scales, as indicated by the wide $q^{-1}$ scaling regime in Fig. 1, red.
Arrested turbulence
We now study the case $S\nu<-1$, realized with extensile flow-aligning rods ($S=+1$ and $\nu=-1.1$). While the system still evolves chaotically, large-scale flows are much weaker than in the other cases, with the $q^{-1}$ power law spanning a narrower range of scales (Fig. 1, blue). Rather than large-scale jets, the dominant flows are at shorter length scales, as indicated by the enhanced peak of the velocity power spectrum (Fig. 1, blue). These smaller-scale flows form streams that are strongly localized along the bending domain walls where the active stress gradients are larger (Figs. 1 to 1).
In contrast to the non-arrested case, here a connected wall pattern extends across the entire system (Fig. 1). The Frank free energy spectrum is more narrowly peaked at the characteristic length scale (Fig. 1, blue), compared to the non-arrested case (Fig. 1, red). The same peak is also prominent in the nematic tensor power spectrum (Supplementary Fig. 2, blue). Strikingly, the nematic wall pattern becomes far less dynamic; the walls and their associated streams become locked in a tree-like pattern. We term this phenomenon dynamical arrest. The weak large-scale chaotic flows make the pattern fluctuate only slightly (Supplementary Movies 2 and 4, right). The dynamical arrest manifests in much longer correlation times, especially of the nematic tensor field at large scales (Fig. 1, blue). The tree-like pattern of domain walls is characterized by unique connectivity properties, which we elaborate on later. We conclude that, in the absence of topological defects, the combined effects of extensile stress and rod-like flow-alignment give rise to the arrest of active nematic turbulence.
[…]
Figure 2: Wall coarsening, branching, and dynamical arrest. Sequential snapshots from a single simulation in the extensile flow-aligning regime, showing the coarsening of bending walls (2, top row), the zigzag instability of the coarsened configuration, followed by wall branching and rapid tip growth (2, second row), and the evolution towards the arrested tree pattern (2, bottom row). The initial condition was set to a striped pattern of wavelength that matches the typical length selected in the chaotic regime (see first and final panels). A perturbation along the $y$-axis triggers the coarsening of one-dimensional stripes, confirming that straight anti-parallel walls do not have a preferred separation. The length selected in the tree-like pattern is determined by two-dimensional interactions, with growing branches avoiding preexisting walls. Parameter values are: $R=1$, $S=1$, $\nu=-1.1$, and $A=19692$ (chosen so that the system size roughly equals six times the selected wavelength).
Wall dynamics and mechanism of arrest
To understand the mechanism of arrest, we note that the spontaneous-flow instability, which ultimately drives turbulence, generates shear flow 46. This shear produces antiparallel flows that concentrate along the nematic walls46; 41; 42; 47; 33; 37. The walls are directed lines, with their direction defined by the vector $(\partial_{y}\theta,-\partial_{x}\theta)$, which is perpendicular to the gradient of the angle and coincides with the flow direction in the arrested regime.
Quasi one-dimensional wall patterns tend to coarsen, leading to larger domains (Fig. 2 and Supplementary Movie 4, right). Since the director orientation is roughly uniform within such domains, the same instability prompts distortions in a perpendicular direction33; 47. This process induces sequential ‘zig-zag’ instabilities, leading to wall branching and progressively smaller domains, until a characteristic wavelength is selected (Fig. 2 and Supplementary Movie 4, right) 19; 33. When an open end of a branch approaches another wall perpendicularly, it is deflected and becomes antiparallel to that wall. The growth of branches typically stops when they are blocked by a preexisting branch point. Ultimately, these dynamics lead to a grid-locked pattern (Fig. 2 and Supplementary Movie 4, right).
When a growing wall is trapped by a branch point, the outflowing stream at the end point of the wall is diverted to both sides so that material can join the antiparallel streams of neighbouring walls (see Fig. 1, around the black circles). In other words, a stagnation point is formed as flows arrive in the direction of the growing wall and leave in the perpendicular direction, to which rods tends to align. For bending walls—induced by an extensile stress—the director is perpendicular to the wall and therefore aligned with the extensional axis of the flow. We thus postulate that the structure around the stagnation point acts as a trap, and its stability is key to understanding how flow alignment promotes the arrest of active turbulence.
[…]
Figure 3: Domain wall nodes and their pseudo-topology. In all diagrams, black directed lines represent nematic bending walls, with gray lines tracing the director and cyan indicating the bending energy. 33, Nodes are either startpoints (green, negative pseudo-defects), branchpoints (blue, negative pseudo-defects), or endpoints (red, positive pseudo-defects). The orange curves illustrate the paths excluding walls used to define the pseudo-charge (Supplementary Note). 33, Two ways in which a startpoint-endpoint pair of opposite charge can be born. The inverse processes of (3) (complete wall dissolution) and (3) (local wall completion) result in the annihilation of such pairs. 3, Wall branching gives birth to a branchpoint-endpoint pair, also of opposite charge. A connected pair may annihilate via branch retraction. 3, A branchpoint (3) transitions into a startpoint when one outgoing wall disjoins. The inverse process corresponds to a startpoint (3) joining with a bare wall. 3, Connected branchpoints can shrink their connecting branch, transiently creating a $-2Q$ pseudo-charged structure, exchange outgoing walls and extend again in the perpendicular direction.
The nodes of nematic domain walls define pseudo-defects
The domain walls emerging in our system are typically of a single type (either splaying or bending, see Figs. 1 and 1). Such walls generically feature three types of nodes:
  1. 1.
    Startpoints, with one outgoing wall and no incoming wall (Fig. 3),
  2. 2.
    Branchpoints, with one incoming wall and two outgoing walls (Fig. 3),
  3. 3.
    Endpoints, with one incoming wall and no outgoing wall (Fig. 3).
A node with two incoming walls and one outgoing wall is only possible by combining different wall types (Supplementary Fig. 3). We further note that startpoints are very common in the non-arrested (contractile aligning) regime but become increasingly rare over time in the arrested (extensile aligning) regime.
We define these three types of points as pseudo-defects. Unlike proper topological defects, they are not phase singularities of the director field, and they do not carry a topological charge. Nevertheless, the director varies around them in a manner that resembles true nematic defects, except for its sharp counter variation across the walls. We thus assign the nodes a pseudo-charge, defined as the number of turns of the director along a path that surrounds the pseudo-defect, excluding the sharp variation across all intersecting walls (Figs. 3, 3 and 3). The excluded walls in this definition are analogous to the so-called Dirac strings, which are constructs that enable magnetic monopoles in Maxwell’s equations48 (see discussion in Supplementary Note). With this definition, endpoints have a positive pseudo-charge $+Q$ (Fig. 3) and both startpoints and branchpoints have pseudo-charge $-Q$ (Figs. 3 and 3). The pseudo-charge depends on $\nu$. For $S\nu>-1$, we have $Q\simeq 1/2$, while for $S\nu\leq-1$, we have $Q\simeq\frac{1}{2}\left[1-\frac{1}{\pi}\arccos{(|\nu|^{-1}})\right]$ (see details in Supplementary Note). The total pseudo-charge contained in a closed domain is $mQ$ where $m$ is the difference between the number of walls crossing the path inwards and outwards (Supplementary Note). All pseudo defects have polarity, defined by the direction of the outgoing wall for startpoints (Fig. 3) and the direction of the incoming wall for branchpoints and endpoints (Figs. 3 and 3).
The topological pseudo-charge is preserved by the dynamics of the director. Pseudo-defects with opposite charges are created in pairs and annihilated in pairs (Figs. 3, 3 and 3), similarly to actual nematic defects. A wall endpoint cannot merge with a preexisting wall. Instead, it can connect with a startpoint ahead of it (the inverse process of Fig. 3) or, through wall shrinkage, it can annihilate with the nearest negative pseudo-defect it is connected to (the inverse processes of Figs. 3 and 3). Notably, this nearest connected point does not need to be its birth pair and may be quite a long distance away, as we often observe in the arrested regime. In addition to pair birth and annihilation, negative pseudo-defects can transition from one to the other while conserving the pseudo-charge (Fig. 3 and its inverse process). In another pseudo-topological transition, connected branchpoints can exchange branches (Fig. 3). This process is akin to a T1 transition in foams and tissues, except that our lines are directed, and branching nodes are restricted to having one incoming wall and two outgoing walls. As a result, T2 transitions are not possible.
One can imagine other transitions between wall nodes that are permitted by the pseudo-topological rules. To the best of our understanding, such processes can be viewed as either analogues or sequences of the basic transitions shown in Fig. 3. Moreover, all processes involving startpoints are rare in arrested wall networks.
[…]
Figure 4: Motifs of arrested wall networks and experimental snapshots of microtubule-kinesin active nematics. In all diagrams, lines, nodes and colors are as defined in Fig. 3. 44, Basic network motifs. The anchoring motif (4) is made of an endpoint and branchpoint that meet head-on, with the endpoint trapped between the two outgoing walls of the branchpoint. In motif (4) the endpoint meets the branchpoint from one of its sides, i.e. between the incoming wall and an outgoing wall. The dashed line traces a weak distortion, indicating that the wall associated with the endpoint tends to align its direction with the outgoing wall on the opposite side of the branchpoint. The motifs (4,4) involve a single pseudo-defect interacting with a bare wall. These, along with (4), do not follow the tendency to have strictly antiparallel walls. 4, Composite motif schematic (left) and one formed spontaneously in a simulation (right). The stream plot on the right represents the flow, with black indicating maximal $|\bm{v}|$ and full transparency indicating $|\bm{v}|=0$. The gray background is the line-integral-convolution representation of the nematic director $\bm{n}$. Parameter values and color legend for splay and bend distortions are as in Fig. 1. 44, Raw fluorescence images from experiments (left panels) and overlaid schematic drawings (right panels) depicting domain walls, pseudo-defects and actual ±$\pm 1/2$ defects in white. 4, Taken from a movie in 49 (courtesy of Guillaume Duclos), which shows the evolution of the microtubule-based nematic following the bending instability of the aligned state. 4, Taken from a movie in 50 (courtesy of Pau Guillamat), which shows a turbulent transient with all types of pseudo-defects and actual nematic defects. Note how walls may also originate from true $+1/2$ defects and be absorbed by true $-1/2$ defects.
Motifs of arrested wall networks
In the arrested turbulence regime (with $S\nu<0$), the processes of wall folding, branching, extension, and arrest produce a directed tree-like network of self-avoiding walls (Fig. 2). This network does not form closed loops in a simply-connected domain because such a loop would enclose a topological charge of 1, implying actual nematic defects, which our system forbids.
We find that endpoints propagate rapidly in the direction of their polarity, until they are blocked by other walls. Startpoints, on the other hand, move quickly in the opposite direction of their polarity until their wall dissolves or they merge with another wall, transitioning into a branchpoint. Unlike endpoints and startpoints, branchpoints do not move in the direction parallel or antiparallel to their polarity. Rather, when they are created (Fig. 3), they tend to deflect only slightly in the direction of the newly formed branch (Fig. 2). Hence, in the sense of their mobility, endpoints are akin to actual $+1/2$ nematic defects and branchpoints are similar to $-1/2$ defects, which are not mobile.
Ultimately, the pseudo-defects—primarily endpoints and branchpoints—maintain a steady distance from neighbouring walls and organize into motifs like those shown in Figs. 4 to 4. A frequent motif, which we call the anchoring motif (Fig. 4), is the one discussed previously in relation to the stagnation points in Fig. 1. We postulate that this structure might be particularly stable because neighbouring walls are anti-parallel, as promoted by the spontaneous-flow instability. In contrast, the other basic motifs (Figs. 4 to 4) feature some neighbouring walls that are parallel. For smaller system sizes, closer to the transition to chaos, we observe that only the anchoring motifs survive in the long time limit. For large system sizes, a variety of composite structures combining all types of motifs can emerge and persist for long times. One such structure is shown in Fig. 4.
We identified bending domain walls, pseudo-defects and our motifs in images and videos of microtubule-based active nematics 51; 50; 15; 49; 52 (Figs. 4 to 4). In Fig. 4, the spontaneous-flow instability leads to the growth of bending domain walls and the emergence of pseudo-defects. These tend to organize in interlaced anchoring motifs, which are also frequently observed in our simulations. In Fig. 4, pseudo-defects are seen to coexist with actual nematic defects (white). Since the latter may also act as origin points or termination points of walls, there may be an interesting interplay between both types of structures in organising the dynamics of the system. It is important to note that in these experimental examples the system does not maintain an arrested state. We expand on this aspect further in the final discussion.
[…]
Figure 5: Ageing of the arrested wall network. 5,5, Skeleton of the domain walls (black) with startpoints, branchpoints and endpoints (green, blue and red triangular nodes) at an early time (5) and a late time (5). The detection of the network skeleton and its nodes is described in Methods .3, Supplementary Fig. 4. 5, Evolution of the number of startpoints (green), branchpoints (blue) and endpoints (red). In the initial transient, sequential ‘zig-zag’ instabilities result in the proliferation of both branchpoints and endpoints. Once the wall pattern establishes a wavelength, the system ages slowly as some endpoints retract and annihilate with their connected branchpoint, while others extend (Supplementary Movie 5). Throughout the simulation, there are frequent transitions between branchpoints and startpoints, though the number of startpoints remains low. Additionally, the detection algorithm is not perfect, occasionally misidentifying endpoints or branchpoints as startpoints and vice versa. Parameter values were set to $R=1$, $S=1$, $\nu=-0.9$, and $A=3.2\times 10^{5}$
Ageing dynamics
The dynamics of active nematics are not relaxational, and the arrested patterns reported here co-exist with a background of weak large-scale chaotic flows. Nevertheless, these dynamics share interesting similarities with relaxational systems that exhibit frustration. In particular, trapping motifs (Figs. 4 to 4) might promote arrest similarly to how cages and locally-favoured structures promote arrest in gels and glasses 53; 54; 55. Additionally, the wall network undergoes a slow relaxation that is reminiscent of the ageing of gels and glasses. The background turbulent-like flows produce small low-$q$ fluctuations that slowly evolve the pattern. This includes jittering the walls and inducing some pseudo-topological transitions in the network; primarily those depicted in Figs. 3, 3 and 3 and their inverse processes. Crucially, these represent a highly restrictive subset of transitions compared to the full range of processes enabled when startpoints are more prominently present. Through this fluctuating process, we find that the number of pseudo-defects gradually decreases (Fig. 5, Supplementary Movie 5), mostly via annihilation of connected branchpoint-endpoint pairs. As some walls shrink and disappear from the network, others extend, thereby maintaining the selected wavelength. At long times ($\sim 10^{3}\,\tau_{\text{a}}$), the evolution slows down and the number of pseudo-defects seems to approach a steady state (Fig. 5).
[…]
Figure 6: Unicursal labyrinths. 6, Two examples of arrested tree-like patterns formed by domain walls, with $S\nu=-0.9$ (left) and $S\nu=-1.1$ (right). 6, Processed negatives of (6) showing how domain walls (black) enclose a small and even number of unicursal labyrinths (4 connected regions on the left, 2 on the right) that span the entire system (Methods .3, Supplementary Fig. 5). The walls separating different labyrinths are self-connected across the periodic boundaries, unlike all other branches which feature endpoints. 6, The mythical labyrinth of Crete, shown on a silver coin from around 400 BC (by AlMare, reproduced from Wikimedia Commons, creative commons license CC BY-SA 3.0). 6, The labyrinth in the Chartres cathedral (courtesy of Père Emmanuel Blondeau and the secretariat de la Cathédrale Notre-Dame de Chartres). 6, Meander pattern known as the Greek key encircling a box (pyxis) from around 850 BC (reproduced from the Metropolitan Museum of Art). 6, Out-of-plain magnetization domains of a ferrimagnetic garnet film, with black and white regions indicating opposite magnetizations. The amplitude of an external magnetic field favouring the black domain is gradually reduced from left to right, resulting in growth by folding of the white unicursal domain. Frames are desaturated snapshots from a video courtesy of Pietro Tierno.
The tree-like wall network encloses unicursal labyrinths
When there are no startpoints, the gaps between the walls form unicursal labyrinths of a type known as meanders, defined by having a single path with neither bifurcations nor dead ends. A typical wall network encloses a small and even number of such labyrinths, with each one closing on itself through the periodic boundaries, i.e., through the hole of the torus, which is not a simply-connected domain (Figs. 6 and 6). On the other hand, the wall network bounding these unicursal paths forms a maze, defined as having multiple bifurcations and dead ends, corresponding to our branchpoints and endpoints. The specific walls separating the different unicursal labyrinths are initially generated by the spontaneous-flow instability. There are as many of these primary walls as there are unicursal labyrinths, and they also close on themselves through the periodic boundaries. Since primary walls are the only paths on the network that avoid dead ends, they can be thought of as the solutions to the maze.
Paintings, rock-carvings, ornaments and monuments depicting unicursal labyrinths predate recorded history and have remained common in artistic and spiritual designs 56. Renowned examples include the Cretan labyrinth of Greek mythology (Fig. 6), and the labyrinth decorating the floor of Chartres cathedral (Fig. 6). However, unlike ours, these and most other Labyrinth designs feature a dead end in the center. On the other hand, meanders (Fig. 6) may close on themselves through the hole of the torus. In pattern-forming systems, unicursal labyrinths are rare. They were obtained under special conditions in simulations of a modified Swift-Hohenberg equation 57, and in experiments with ferrimagnetic garnet films in specific experimental protocols 58; 59; 60 (Fig. 6). In our case, they result from an emerging topology of domain wall networks and they occur spontaneously and generically in the arrested regime.
Discussion
In this work, we showed that flow alignment—the tendency of liquid crystals to reorient under shear—dramatically influences the spatiotemporal structure of defect-free active nematic turbulence (Fig. 1). For contractile flow-aligning active nematics, the chaos is characterized by fragmented nematic patterns, reduced total Frank free energy, and strong self-similar flows across a broad range of scales. We associate this strong large-scale turbulence to the underlying metastability of uniform and wall-laden states 37. On the other hand, despite their chaotic dynamics, we find that extensile aligning active nematics self-organize into an arrested tree-like pattern of nematic domain walls, with the flow strongly channeled along them (Figs. 1 and 1). This finding reveals a mechanism of pattern formation mediated by active turbulence, complementing those found in other studies 61; 62; 63. In particular, our patterns enclose unicursal labyrinths (Fig. 6), which do not appear spontaneously in other pattern-forming systems.
The labyrinths in Fig. 6 span the entire system, but how far can they possibly extend? Further increasing the system size widens the range of chaotic self-similar flows, identified by the $q^{-1}$ scaling of the velocity power spectrum (Fig. 1, blue). These flows become stronger at larger scales, and hence they could potentially disrupt the arrested wall pattern for sufficiently large systems. At the scale where chaotic large-scale flows become as strong as the wall streams, which sustain the domain walls, the tree-like network could lose its connectivity and fluctuate more strongly. In fact, for the system size studied here, this competition between wall streams and chaotic large-scale flows is observed for $\nu=0$ (Fig. 1, black). In this case, we observe a wall network, locally similar to the arrested ($S\nu=-1.1$) pattern, but sporadically fractured in random positions (Supplementary Fig. 1, Supplementary Movie 3). In line with our hypothesis, the fractured areas where the wall pattern dissolves (dark regions in the Frank free energy density) appear to coincide with hot spots of large-scale flows. The wall pattern is subsequently restored at fractured regions because large jets are not persistent, as shown by the low correlation times of low-$q$ flow in Fig. 1, dashed black.
The arrested state does not show long-range correlations in the nematic order parameter (Supplementary Fig. 2, blue). This is reminiscent of some passive systems, such as spin ice, that feature frustration in the absence of long-range geometrical correlations64. Similarly, our pattern features short-range disorder and is yet connected over long distances. Therefore, we propose that the unicursal labyrinths exhibit a form of long-range topological order stemming from emergent connectivity rules associated with the folding, branching, extension and self-avoidance of the domain walls.
In the fully-developed turbulent state, both the arrested and non-arrested regimes manifest a selected wavelength. This wavelength is significantly greater than the active length $l_{\text{a}}$ (Fig. 1), and it also exceeds the critical wavelength of the spontaneous-flow instability. The wavelength selection mechanism is inherently nonlinear and two-dimensional; it is based on a balance between the coarsening of local striped patterns, which tends to enlarge the characteristic wavelength (Fig. 2), and the bending and folding of walls that tends to reduce it (Fig. 2). A deeper understaning of the selection mechanism and its dependence on parameters remains an open question.
Whereas our model is defect-free, most experiments and simulations in active nematics have been conducted in defect-laden regimes. Yet, the physics of the two regimes may intersect, perhaps similarly to the crossovers found in polar fluids 65; 66. Actual nematic defects tend to nucleate along nematic domain walls, thereby dissolving them 43; 44; 45; 34. Additionally, nematic constituents such as microtubules are depleted at defects (Fig. 4), which lowers their energy cost and favours their nucleation67. The frequent dissolution of domain walls by defects explains why, in the instances where our motifs can be seen (Figs. 4 to 4), they do not last for long. We therefore propose that, in addition to their well-known role as drivers of turbulence by generating vorticity 32, defects also play an important role as enablers of stronger turbulent flows by breaking the potential grid-lock, i.e., preventing dynamical arrest.
Directly testing our predictions for defect-free active nematics presents an experimental challenge. This would require searching for active materials for which the nucleation of defect pairs were significantly more costly than in the experimental realizations studied thus far. Defect-free active turbulence should be achievable for sufficiently large system sizes $L$, ensuring large dimensionless activity $A=(L/\ell_{\text{a}})^{2}\gg 1$, while keeping the defect-core size $a$ small enough ($\ell_{\text{a}}\gg a$) to preclude defect nucleation.
In conclusion, our study of defect-free active nematic turbulence highlights the crucial role of topology beyond the existence of topological defects. We argue that the connectivity of domain walls, the existence of pseudo-defects, and their interactions are key to understanding the dynamics of active nematics, from strong large-scale turbulence to dynamical arrest. Furthermore, when true defects are present, their interplay with pseudo-defects likely adds another intriguing layer of complexity. Our results open new avenues for experimental analysis and design, particularly in tracing the intricate web of domain walls, their orientation, and their nodes.
Acknowledgments
Funding from the Spanish Ministry for Science and Innovation MICCINN/FEDER (PID2022-137713NB-C22 to IL and JC) and the Generalitat de Catalunya (AGAUR SGR-2021-00450 to IL and JC and ‘ICREA Academia’ award to JC) is acknowledged. IL acknowledges support from the Flatiron Instutute, part of the Simons Foundation. RA thanks the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the Programme New Statistical Physics in Living Matter, supported by EPSRC Grant Number EP/R014604/1, when work on this paper was undertaken.
Author contributions
IL performed all the simulations and analyses. JC conceived and supervised the research. IL and JC developed the theory. All authors discussed the results and wrote the manuscript.
Competing interests
The authors declare no competing interests.
Data and code availability
The data and code supporting the findings of this study are available from the corresponding author, IL, upon reasonable request.
Supplementary Figures
[…]
Supplementary Figure 1: Active nematic turbulence in the absence of flow-alignment. Snapshots from a simulation of defect-free active nematic turbulence with parameter values $R=1$, $\nu=0$, and $A=3.2\times 10^{5}$. We chose a contractile stress ($S=-1$), but the results are exactly equivalent in the extensile case ($S=+1$) when the director is rotated ($\theta\to\theta\pm\pi/2$). 1, The flow field; black curves are streamlines, and the color indicates the speed (Supplementary Movie 3, left). 1, Frank free energy density $\sim|\nabla\theta|^{2}$, with high-intensity lines corresponding to nematic domain walls (Supplementary Movie 3, right). Note how sparse peaks in the flow (red regions in 1) tend to coincide with perturbed wall patterns (dark regions in 1). 11, Zooms in the contractile (1) and extensile (1) cases highlighting the type of nematic distortion as well as the interplay between nematic walls and flows. The gray-scale background is the line integral convolution representation of the director field $\bm{n}$. Magenta and cyan intensities respectively represent splay $(\nabla\cdot\bm{n})^{2}$ and bend |×$|\nabla\times\bm{n}|^{2}$ contributions to the Frank energy density. The black arrows represent the flow field $\bm{v}$.
Supplementary Figure 2: Power spectrum of the order parameter in active nematic turbulence. Shown is the power spectrum of the nematic tensor, $\hat{q}_{\alpha\beta}=n_{\alpha}n_{\beta}-1/2\,\delta_{\alpha\beta}$, in the contractile flow-aligning ($S\nu=+1.1$, red), extensile flow-aligning ($S\nu=-1.1$, blue), and non-aligning ($\nu=0$, black) cases. For the definition of $W(q)$ see Methods .2. Lines represent a smoothed (Gaussian) interpolation of the computed data points. Parameter values are as in Fig. 1 and Supplementary Fig. 1. For arrested turbulence (blue), this spectrum features a narrow peak with its associated harmonics. For strong large-scale turbulence (red), such peaks are washed out. In all regimes, low-$q$ correlations are essentially absent.
[…]
Supplementary Figure 3: Domain wall nodes in defect-free active nematics (expanded topology). 33, When walls are dominated by one type of distortion (either bend or splay), the possible nodes are startpoints (green, negative pseudo-defects), branchpoints (blue, negative pseudo-defects), and endpoints (red, positive pseudo-defects). 3, Mergepoints (yellow, positive pseudo-defects) feature two incoming walls of the same type and one outgoing wall of the other type. While such combinations are topologically possible, they are not typically featured in our simulations.
[…]
Supplementary Figure 4: Skeleton detection and pseudo-defect classification. 4, Grayscale plot of the Frank free energy density, with black corresponding to $\left|\nabla\theta\right|^{2}=0$ and white corresponding to $\left|\nabla\theta\right|^{2}\geq 0.5\,\text{max}\left|\nabla\theta\right|^{2}$ to facilitate binarization. 4, Periodic expansion of (4) to avoid false detection of nodes along the original square boundaries (white outline). 4, Local adaptive binarization of (4), followed by removal of small fragmented components not counted as walls. 4, Recursive thinning of (4), then a step of dilation to fill up tiny gaps within some walls, an additional thinning step to recover the skeleton, and a pruning step to discard open branches shorter than a small threshold. 4, Using the MorphologicalTransform tool we detect and highlight the skeleton nodes. 4, We select only those nodes confined to the original periodic square. 4, For the SkeletonEndPoints, we compute the vectors pointing outwards of their associated branch (black arrows) by finding the nearest neighbouring point on the wall skeleton. 4, The wall polarity associated with each SkeletonEndPoint is shown in red (green) if its scalar product with (4) is positive (negative). 4, Branchpoints (blue nodes), endpoints (red nodes) and startpoints (green nodes) are plotted with their orientation given by the local wall polarity.
[…]
Supplementary Figure 5: Labyrinths detection in arrested wall networks. 5, Grayscale plot of the Frank free energy density, with white corresponding to $\left|\nabla\theta\right|^{2}=0$ and black corresponding to $\left|\nabla\theta\right|^{2}\geq 0.5\,\text{max}\left|\nabla\theta\right|^{2}$ to facilitate binarization. 5, Periodic expansion of (5) to capture connected domains that meander through the periodic square boundaries. 5, Local adaptive binarization of (5), followed by a filling operator to discard isolated black fragments. 5, Finding the connected morphological domains in (5). 5, Cropping out the periodic square domain from (5). 5, Applying our algorithm for relabeling domains that are connected through the left-right and top-bottom boundaries.
[…]
Supplementary Figure 6: Verification of the numerical scheme. 6, Consecutive snapshots from a simulation initiated with a small bending modulation about the uniformly aligned state: $\theta(\bm{r},t=0)=\epsilon\cos(2\pi x)$, with $\epsilon=0.01$. Parameter values were set to $R=1$, $\nu=-1.1$, $A=1000$, and $S=1$ (extensile stress). Numerical parameters were set to $N=64$, $dt=0.01$, and $D=0$ (zero noise). Color indicates the Frank energy, $|\nabla\theta|^{2}$, and white lines trace the director $\bm{n}$. 6, We plot the numerical data corresponding to a horizontal slice, $\theta(x,y=0)$, for increasing times, starting from $t=0$ (lightest gray) up to $t=40\tau_{\text{a}}$ (black) in intervals of $2.5\tau_{\text{a}}$. It is shown that the angle profile relaxes on the predicted 1D steady state (dashed blue), obtained for the same parameter values as explained in our recent work37. 6, The maximum angle is plotted as function of time in our simulation (black). First, it is shown that the modulation grows exponentially in time with the growth rate $\Omega$ matching the linear stability dispersion relation 37 (cyan). The gray inset is a log-scale plot comparing the simulation results with this explicit prediction for small perturbations. Additionally, it is demonstrated that the maximal angle in the simulation relaxes at long times to the value predicted by the 1D steady state 37 (dashed blue).
Supplementary Movie Captions
Supplementary Movie 1 Strong large-scale turbulence in the contractile aligning regime. The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Figs. 1 and 1, which are snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the Movie is $2500\tau_{\text{a}}$.
Supplementary Movie 2 Arrested turbulence in the extensile aligning regime. The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Figs. 1 and 1, which are snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the Movie is $2500\tau_{\text{a}}$.
Supplementary Movie 3 Active turbulence in the absence of flow alignment (tumbling regime). The left panel shows the evolution of the flow field $\bm{v}$, with color indicating speed and black arrows representing streamlines. The right panel depicts the evolution of the Frank free energy density, $|\nabla\theta|^{2}$. The parameters are the same as in Extended Data Fig. 1a and 1b, which are snapshots from the same simulation. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the Movie is $2500\tau_{\text{a}}$.
Supplementary Movie 4 Evolution from stripes to turbulence (contractile aligning vs extensile aligning). The evolution of the Frank free energy density, $|\nabla\theta|^{2}$, in the contractile aligning case ($S\nu=+1.1$, left) and the extensile aligning case ($S\nu=-1.1$, right). The wavelength of the initial stripes—with splay walls on the left and bend walls on the right—was chosen to match the selected wavelength in the fully-developed turbulent state. In both cases, the stripes coarsen and then experience a ‘zigzag’ instability that folds the walls. In the contractile case (left), folded walls tend to dissolve and reappear perpetually. In the extensile case (right, Fig. 2), walls are more robust and tend to branch. As growing branches avoid other walls, the system reaches a grid-locked state that fluctuates only slightly. Parameters were set to $R=1$, $\nu=-1.1$, and $A=19692$ (chosen so that the system size roughly equals 6 times the selected wavelength). The interval between movie frames is 50 computational time iterations, corresponding to a time interval of $1\tau_{\text{a}}$. The total duration of the Movie is $540\tau_{\text{a}}$.
Supplementary Movie 5 Evolution of the wall-network skeleton and its nodes. Skeleton of the domain walls (black) with startpoints, branchpoints and endpoints shown as green, blue and red triangular nodes, respectively (see Fig. 3). The detection of these structures is described in Methods C and Extended Data Fig. 5. This movie demonstrates the pseudo-topological transitions mediating the ageing of an arrested wall network. Parameter values are as in Fig. 5. The interval between movie frames is 500 computational time iterations, corresponding to a time interval of $5\tau_{\text{a}}$. The total duration of the Movie is $1715\tau_{\text{a}}$.
Supplementary Note:
Topological pseudo-charge of domain wall nodes
In arrested and nearly arrested states, the domain walls are organized in a tree-like network. One can associate a planar, directed graph to this network, with three types of nodes: startpoints, endpoints, and branchpoints, as described in Figs. 3 to 3. In the long-time asymptotics, the startpoints are only residual, and the ageing tree is mostly formed by endpoints and branchpoints (Fig. 5, Supplementary Movie 5). While these nodes are not topological defects of the nematic field, one can assign to them a pseudo-charge which is conserved and contains useful information of topological nature, as discussed in this section. We emphasize that this pseudo-topology is an effective emergent property of the dynamics, that holds only for the appropriate parameter regimes and time asymptotics. Therefore, pseudo-defects are dynamically stabilized but are not topologically stable in the sense that they can be removed by a continuous perturbation of the nematic field.
To define the topological pseudo-charge, we first consider the change of angle across a nematic wall. To this end, we look for steady-state solutions of the director field in 1D, that is, with translational invariance in the direction along the wall. In this case, the angle varies only along the transverse coordinate $s$. In the steady state, the director profile across the wall obeys the 1D stationary equation 37
$\ell_{\text{a}}^{2}\partial_{s}^{2}\theta=\frac{S(1+\nu\cos 2\theta)(\sin 2\theta-\sin 2\bar{\theta})}{4+R+R\nu^{2}+2R\nu\cos 2\theta},$ (S1)
where $\bar{\theta}$ is the orientation of the nematic director at the center of the domain wall. This equation supports sigmoidal angle profiles with an interface of thickness of the order of the active length $\ell_{a}$ that continuously connects two plateaux of uniform angle. As the interfaces separate regions of uniform angle, they are referred to as domain walls, also called kinks and antikinks in other contexts.
For $S\nu>-1$, the two plateaux values of the nematic angle are $\theta=\bar{\theta}\pm\frac{\pi}{2}$, so the total rotation of the nematic across the wall is $\Delta\theta=\pm\pi$. For $S\nu<-1$, the plateaux angles correspond to the so-called Leslie angle defined by $01+\nu\cos 2\theta_{L}=$ 0, which amounts to an angle rotation of $\Delta\theta=\pm(\pi-\arccos{|\nu|^{-1}})$ ) across the wall. We remark that the total angle variation across domain walls is independent of both the active length $\ell_{\text{a}}$ and the viscosity ratio $R$. For a domain wall centered at $s=s_{i}$, we define the smooth step functions $\Theta^{\pm}(s-s_{i})$ that connect the plateaux values via a total angle increment $\Delta\theta$, which is positive for $\Theta^{+}$ and negative for $\Theta^{-}$. The angle varies at a scale $\sim\ell_{\text{a}}$ about $s_{i}$. Similarly, the derivative with respect to the spatial coordinate $s$ defines two peaked functions $\dot{\Theta}^{\pm}(s-s_{i})$ such that, in the sharp-wall limit,
$\lim_{\ell_{\text{a}}\rightarrow 0}\dot{\Theta}^{\pm}(s-s_{i})=\Delta\theta\,\delta(s-s_{i})\equiv\pm\frac{Q(\nu)}{2\pi}\delta(s-s_{i}).$ (S2)
Here, $Q(\nu)$ is the elementary pseudo-charge function, defined as
[…]
which gives, for $S\nu<0$,
$Q(\nu)=\frac{1}{2}\left(1-\frac{1}{\pi}{\cal H}(|\nu|-1)\arccos{|\nu|^{-1}}\right)>0,$ (S4)
where ${\cal H}$ is the Heaviside step function. The elementary pseudo-charge function $Q(\nu)$ takes the value $1/2$ in the tumbling regime, and it spans the interval from $1/2$ to $1/4$ as $|\nu|$ is increased beyond $1$. The functional form of $Q(\nu)$ is plotted in Fig. S1.
To define a topological pseudo-charge of the pseudo-defects, and of any closed domain in general, we recall that the total angle variation along a closed loop must be
[…]
Consider, for instance, an isolated endpoint pseudo-defect (Fig. 3), formed by a wall that points towards it, and consider a counterclockwise closed path $\Gamma$ that contains it and intersects the center of the wall at $s=s_{0}$. We define the pseudo-charge of this pseudo-defect as
[…]
which coincides strictly with $Q(\nu)$ in the sharp-wall limit. To subtract the localized fast variation of $\theta$ across the wall we need to write $\Theta^{-}(s-s_{0})$ because the nematic is rotating clockwise across the wall (decreasing angle), if the wall is oriented towards the interior of the path that encloses the endpoint. In general, the sense of rotation of the nematic across a wall is opposite (equal) to the sense of rotation of the closed loop when the wall points towards the interior (exterior) of the loop.
Figure S1: $Q$ as a function of $|\nu|$. Functional form of the elementary charge $Q(\nu)$. Positive pseudo-defects have charge $+Q$, and negative ones $-Q$. For the tumbling regime $Q=1/2$ and for the aligning regime $Q$ decreases monotonically from 1/2 to 1/4.
Similarly, for an isolated branchpoint we define the pseudo-charge as the rotation of the director in regions excluding the walls along a closed loop enclosing the branchpoint (Fig. 3). In this case, we must subtract the fast variation across the three walls intersected by the integration path. Since there is always one wall pointing towards the interior and two pointing towards the exterior, the pseudo-charge for a branchpoint, in the sharp-wall limit, will be $-Q(\nu)$.
For a general configuration with an arbitrary number of pseudo-defects and associated walls, the total pseudo-charge in the interior of a closed loop will given by the sum of the pseudo-charge of all the pseudo-defects in it. We remark that it is not necessary to know the detailed structure of the nematic field in the interior, but only in the contour, as in the definition of the actual topological charge. In fact, the total pseudo-charge in the interior of an arbitrary closed loop $\Gamma$ is also given by
$Q_{\Gamma}=(m_{\text{in}}-m_{\text{out}})Q(\nu),$ (S8)
where $m_{\text{in}}$ and $m_{\text{out}}$ are the number of inward and outward domain walls crossed by the contour line, respectively.
In the case of domain walls with a small but finite thickness compared to typical wall separation, the elementary pseudo-charge will slightly differ from the limiting value defined in Eq. S3. In that case, consecutive walls partially overlap so that the angle profile does not reach the asymptotic value of the plateau, and the magnitude of the angle variation across the walls, $\Delta\theta$, is slightly smaller. This is the case in the wall patterns studied in this work, so in the main manuscript we refer to the values taken by the pseudo-topological charge only as $\simeq Q(\nu)$. We remark, however, that the precise value of the elementary charge $Q(\nu)$ is not really informative of the relevant topological structure, that is, how walls are connected, but accounts for the detailed structure of the walls themselves. This may be important to determine their dynamical properties, such as their interactions or their rigidity to deformation. However, it is only the integer counts of positive and negative pseudo-defects (or the counts of inward and outward wall intersections), i.e., the multiples of the elementary charge, that are informative of the overall qualitative structure inside a given domain.
In the sharp-wall limit $\ell_{\text{a}}\rightarrow 0$, the idea of a pseudo-defect is reminiscent of Dirac’s construction of magnetic monopoles, based on attaching to them the so-called Dirac string, which carries a quantized magnetic flux48. In that way, the magnetic monopole can be made consistent with Maxwell’s equations. For instance, the total magnetic flux across a closed surface enclosing the magnetic monopole would be zero, as required by Gauss’ theorem for the absence of net magnetic charge, because the flux generated by the monopole is canceled by that carried out by the string, which also crosses the surface. Similarly, in our problem, the wall attached to the pseudo-defect provides a localized (fast) rotation of the director that cancels the (slow) rotation outside the wall to ensure the lack of true topological charge. In this way, the problem satisfies the local constraint of zero topological charge while mimicking with pseudo-defects the director field around actual topological defects in a large region of space around them.